Marissa has $4.70 in quarters and dimes. If she has 9 more quarters than dimes, how many of each coin does she have?
? dimes
? quarters

Answer by mananth(16946)

dimes x
quarters (x+9)

Value =4.7
10x+25* 1(x+9)=470

10x+25x +225=470

35x+225 =470

x=470 -225
35x=245
/35
x=7 dimes
quarters=16